My question is what is the value of a
a=b+3(4-c)
if. B=5 c=2
what is the answer and can you break it down for me

Answer by stanbon(48558)

My question is what is the value of a
a=b+3(4-c)
if. b=5 c=2
-----------------------
Put the values in place of the letters to get:
---
a = 5 + 3(4 - 2)
---
a = 5 + 3(2)
a = 5 + 6
a = 11
=========================
Cheers,
Stan H.